# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ V-A Uniqueness of Time-Optimal Control
Under Assumption 1, the time-optimal control of problem (1a) is unique in an almost everywhere sense. In other words, if $u=u_{1}^{*}\left(t\right)$ and $u=u_{2}^{*}\left(t\right)$, $t\in\left[0,t_{\mathrm{f}}^{*}\right]$, are both optimal controls of problem (1a), then $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere; hence, $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$.
@@
If $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, then evidently it has $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere since $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u$ and ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{0}$.
@@
Assume that $\exists t^{\prime}\in\left(0,t_{\mathrm{f}}^{*}\right)$, s.t. ${\boldsymbol{x}}_{1}^{*}\left(t^{\prime}\right)\not={\boldsymbol{x}}_{2}^{*}\left(t^{\prime}\right)$. Let $t_{0}=\arg\min\left\{t\in\left(0,t_{\mathrm{f}}^{*}\right):{\boldsymbol{x}}_{1}^{*}\left(t\right)\not={\boldsymbol{x}}_{2}^{*}\left(t\right)\right\}$. Evidently, $0\leq t_{0}<t_{\mathrm{f}}^{*}$. Assume that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)\not={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$. Then, $t_{0}>0$. Due to the continuity of ${\boldsymbol{x}}$, $\forall 0\leq t<t_{0}$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$, which contradicts the assumption. Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ holds. Furthermore, $\forall t\in\left[0,t_{0}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ for $t\in\left[0,t_{0}\right]$ almost everywhere.
@@
According to Bellman’s principle of optimality [20], ${\boldsymbol{x}}={\boldsymbol{x}}_{1}^{*}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}_{2}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ are still the time-optimal trajectories between ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t_{\mathrm{f}}^{*}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{\mathrm{f}}^{*}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$. Let $u_{3}^{*}\left(t\right)=\frac{1}{2}\left(u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)\right)$ and ${\boldsymbol{x}}_{3}^{*}\left(t\right)=\frac{1}{2}\left({\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)\right)$. Evidently, $u_{3}^{*}\left(t\right)$ is feasible for problem (1a) since the feasible set of $\left(u,{\boldsymbol{x}}\right)$ is convex; hence, $u_{3}^{*}\left(t\right)$ is optimal.
@@
Denote that $\forall k=1,2$, the switching law of ${\boldsymbol{x}}={\boldsymbol{x}}_{i}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ is ${\mathcal{L}}_{i}={\mathcal{S}}_{1}^{\left(i\right)}{\mathcal{S}}_{2}^{\left(i\right)}\dots\mathcal{S}_{N_{i}}^{\left(i\right)}$, where $\forall j\in\left[N_{i}\right]$, ${\mathcal{S}}_{j}^{\left(i\right)}=\left(\widehat{{\boldsymbol{A}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{b}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{F}}}_{j}^{\left(i\right)},\widehat{{\boldsymbol{g}}}_{j}^{\left(i\right)},\mathcal{P}_{j}^{\left(i\right)}\right)$ lasts for $t\in\left(t_{j-1}^{\left(i\right)},t_{j}^{\left(i\right)}\right)$. Among them, $t_{0}^{\left(i\right)}=t_{0}$ and $t_{N_{i}}^{\left(i\right)}=t_{\mathrm{f}}^{*}$. Evidently, if ${\mathcal{S}}_{1}^{\left(1\right)}={\mathcal{S}}_{1}^{\left(2\right)}$, then $\forall t\in\left(t_{0},\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}\right)$, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$. Hence, ${\mathcal{S}}_{1}^{\left(1\right)}\not={\mathcal{S}}_{1}^{\left(2\right)}$.
@@
Assume that both ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ are unconstrained arcs. Without loss of generality, assume that $u_{1}^{*}\left(t\right)\equiv{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{\left(1\right)}$; hence, $u_{2}^{*}\left(t\right)\equiv-{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{\left(2\right)}$. Then, $\exists\varepsilon>0$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$ and ${\boldsymbol{C}}{\boldsymbol{x}}_{2}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$. Therefore, ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$. In other words, ${\boldsymbol{x}}_{3}^{*}$ is strictly unconstrained in $\in\left(t_{0},t_{0}+\varepsilon\right)$. According to Proposition 2, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the fact that $u_{3}^{*}\left(t\right)=\frac{u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)}{2}\equiv 0$. Therefore, one of ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ is constrained. Without loss of generality, assume that ${\mathcal{S}}_{1}^{\left(1\right)}$ is constrained.
@@
Assume that ${\mathcal{S}}_{1}^{\left(2\right)}$ is an unconstrained arc. Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}-t_{0}\right)$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, ${\boldsymbol{C}}{\boldsymbol{x}}_{2}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$. Since ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}\leq{\boldsymbol{0}}$, it has ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$. In other words, ${\boldsymbol{x}}_{3}^{*}$ is strictly unconstrained in $\in\left(t_{0},t_{0}+\varepsilon\right)$. According to Proposition 2, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{3}^{*}\left(t\right)=\frac{u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)}{2}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. Note that $u_{1}^{*}\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ and $u_{2}^{*}\left(t\right)\equiv u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{1}^{*}\left(t\right)\equiv u_{0}$. Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ for $\left[t_{0},t_{0}+\varepsilon\right]$, which contradicts the definition of $t_{0}$.
@@
Therefore, both ${\mathcal{S}}_{1}^{\left(1\right)}$ and ${\mathcal{S}}_{1}^{\left(2\right)}$ are constrained arcs. Assume that $\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}\not=\varnothing$. Arbitrarily consider $p\in\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}$. $\forall t\in\left(t_{0},\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}\right)$, $i=1,2$, $\dot{\boldsymbol{x}}_{i}^{*}\left(t\right)=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}_{i}^{*}\left(t\right)$. ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$. Therefore, $\mathcal{P}_{1}^{\left(1\right)}\cap\mathcal{P}_{1}^{\left(2\right)}=\varnothing$. Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{\left(1\right)},t_{1}^{\left(2\right)}\right\}-t_{0}\right)$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $i=1,2$, $p\not\in\mathcal{P}_{1}^{\left(i\right)}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{i}^{*}\left(t\right)+d_{p}<0$. Hence, $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{3}^{*}\left(t\right)+d_{p}={\boldsymbol{c}}_{p}^{\top}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+u{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+d_{p}<0$. According to Proposition 2, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the definition of $t_{0}$. Therefore, $\forall t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ almost everywhere. ∎
@@
Based on Theorem 1, it can be proved that the optimal control of problem (1a) is unique in an almost everywhere sense when there exists at most one chattering limit point. The proof is omitted since it does not contribute to the proposed state-centric necessary condition in Section V-C.
According to Theorem 1, if the time-optimal trajectory is disturbed by a sufficiently small perturbation, then the resulting disturbed trajectory achieves a strictly shorter terminal time than the optimal terminal time. The above observation will be applied in Section V-B.
V-B Feasibility of the Disturbed Bang-Singular-Bang Trajectory

### v2

@@ V-A Uniqueness of Time-Optimal Control
Under Assumption 1, the time-optimal control of problem (1a) is unique in an a.e. sense. In other words, if $u=u_{1}^{*}\left(t\right)$ and $u=u_{2}^{*}\left(t\right)$, $t\in\left[0,t_{\mathrm{f}}^{*}\right]$, are both optimal, then $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ a.e.; hence, $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$.
@@
If $\forall t\in\left[0,t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, then it holds a.e. that $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ since ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{0}$.
@@
Assume that $\exists t^{\prime}\in\left(0,t_{\mathrm{f}}^{*}\right)$, s.t. ${\boldsymbol{x}}_{1}^{*}\left(t^{\prime}\right)\not={\boldsymbol{x}}_{2}^{*}\left(t^{\prime}\right)$. Let $t_{0}=\arg\min\left\{t\in\left(0,t_{\mathrm{f}}^{*}\right):{\boldsymbol{x}}_{1}^{*}\left(t\right)\not={\boldsymbol{x}}_{2}^{*}\left(t\right)\right\}$. Evidently, $0\leq t_{0}<t_{\mathrm{f}}^{*}$. Assume that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)\not={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$. Then, $t_{0}>0$. Due to the continuity of ${\boldsymbol{x}}$, $\forall t\in\left(0,t_{0}\right)$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$, which contradicts the assumption. Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ holds. Furthermore, $\forall t\in\left[0,t_{0}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ implies that $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ for $t\in\left[0,t_{0}\right]$ a.e.
@@
According to Bellman’s principle of optimality [22], ${\boldsymbol{x}}={\boldsymbol{x}}_{1}^{*}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}_{2}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ are still the time-optimal trajectories between ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t_{\mathrm{f}}^{*}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{\mathrm{f}}^{*}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$. Let $u_{3}^{*}\left(t\right)=\frac{1}{2}\left(u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)\right)$ and ${\boldsymbol{x}}_{3}^{*}\left(t\right)=\frac{1}{2}\left({\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)\right)$. Evidently, $u_{3}^{*}\left(t\right)$ is feasible since the feasible set of $\left(u,{\boldsymbol{x}}\right)$ is convex; hence, $u_{3}^{*}\left(t\right)$ is optimal.
@@
Denote that $\forall k=1,2$, the switching law of ${\boldsymbol{x}}={\boldsymbol{x}}_{i}^{*}\left(t\right)$, $t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$ is ${\mathcal{L}}_{i}={\mathcal{S}}_{1}^{(i)}{\mathcal{S}}_{2}^{(i)}\dots\mathcal{S}_{N_{i}}^{(i)}$, where $\forall j\in\left[N_{i}\right]$, ${\mathcal{S}}_{j}^{(i)}=\left(\widehat{{\boldsymbol{A}}}_{j}^{(i)},\widehat{{\boldsymbol{b}}}_{j}^{(i)},{\boldsymbol{F}}_{j}^{(i)},{\boldsymbol{g}}_{j}^{(i)},\mathcal{P}_{j}^{(i)}\right)$ lasts for $t\in\left(t_{j-1}^{(i)},t_{j}^{(i)}\right)$. Among them, $t_{0}^{(i)}=t_{0}$ and $t_{N_{i}}^{(i)}=t_{\mathrm{f}}^{*}$. Evidently, if ${\mathcal{S}}_{1}^{(1)}={\mathcal{S}}_{1}^{(2)}$, then $\forall t\in\left(t_{0},\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}\right)$, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ and ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$. So ${\mathcal{S}}_{1}^{(1)}\not={\mathcal{S}}_{1}^{(2)}$.
@@
Assume that both ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ are unconstrained arcs. Without loss of generality, assume that $u_{1}^{*}\left(t\right)\equiv{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{(1)}$; hence, $u_{2}^{*}\left(t\right)\equiv-{u_{\mathrm{m}}}$ in ${\mathcal{S}}_{1}^{(2)}$. Then, $\exists\varepsilon>0$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$ and ${\boldsymbol{C}}{\boldsymbol{x}}_{2}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$. Therefore, ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$. In other words, ${\boldsymbol{x}}_{3}^{*}$ is strictly unconstrained in $\in\left(t_{0},t_{0}+\varepsilon\right)$. According to Proposition 2, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the fact that $u_{3}^{*}\left(t\right)=\frac{u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)}{2}\equiv 0$. Therefore, one of ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ is constrained. Without loss of generality, assume that ${\mathcal{S}}_{1}^{(1)}$ is constrained.
@@
Assume that ${\mathcal{S}}_{1}^{(2)}$ is an unconstrained arc. Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}-t_{0}\right)$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, ${\boldsymbol{C}}{\boldsymbol{x}}_{2}^{*}\left(t\right)+{\boldsymbol{d}}<{\boldsymbol{0}}$. Since ${\boldsymbol{C}}{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{d}}\leq{\boldsymbol{0}}$, it holds that ${\boldsymbol{C}}{\boldsymbol{x}}_{3}^{*}\left(t\right)+{\boldsymbol{d}}={\boldsymbol{C}}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+{\boldsymbol{d}}<{\boldsymbol{0}}$. In other words, ${\boldsymbol{x}}_{3}^{*}$ is strictly unconstrained in $\in\left(t_{0},t_{0}+\varepsilon\right)$. According to Proposition 2, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{3}^{*}\left(t\right)=\frac{u_{1}^{*}\left(t\right)+u_{2}^{*}\left(t\right)}{2}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. Note that $u_{1}^{*}\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ and $u_{2}^{*}\left(t\right)\equiv u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $u_{1}^{*}\left(t\right)\equiv u_{0}$. Therefore, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$ for $\left[t_{0},t_{0}+\varepsilon\right]$, which contradicts the definition of $t_{0}$.
@@
Therefore, both ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ are constrained arcs. Assume that $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}\not=\varnothing$. Arbitrarily consider $p\in\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}$. $\forall t\in\left(t_{0},\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}\right)$, $i=1,2$, $\dot{\boldsymbol{x}}_{i}^{*}\left(t\right)=\left({\boldsymbol{A}}-{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}_{i}^{*}\left(t\right)$. ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$. Therefore, $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}=\varnothing$. Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}-t_{0}\right)$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $i=1,2$, $p\not\in\mathcal{P}_{1}^{(i)}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{i}^{*}\left(t\right)+d_{p}<0$. Hence, $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{3}^{*}\left(t\right)+d_{p}={\boldsymbol{c}}_{p}^{\top}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+u{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+d_{p}<0$. By Proposition 2, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the definition of $t_{0}$. Therefore, $\forall t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ a.e. ∎
@@
Based on Theorem 1, it can be proved that the optimal control of problem (1a) is unique in an a.e. sense when there exists at most one chattering limit point. The proof is omitted since it does not contribute to the proposed state-centric necessary condition in Section V-C.
According to Theorem 1, if the time-optimal trajectory is perturbed by a sufficiently and finitely small perturbation, then the resulting perturbed trajectory achieves a strictly shorter terminal time than the optimal terminal time. The above observation will be applied in Section V-B.
V-B Feasibility of the Perturbed BBS Trajectory